OpenStudy (anonymous):

Assume that S/A=6.0, D/A=.40, N/E=.80 and A=D+E. Solve for N/S

OpenStudy (anonymous):

Five equations with five unknowns.. lots of ways to do this. Can probably just try different substitutions until you break it all down. A more systematic way would be to put all five equations into a linear equation system or augmented matrix. Have you tried anything so far? Stuck at a particular point?

OpenStudy (anonymous):

No this has got me lost ... I tried to substitute 100 for A and got turned all around.

OpenStudy (anonymous):

Why 100 for A? Just a guess? I checked again, and there's actually only 4 equations for these five unknowns, which means you can't get a unique value for each variable, but the question is asking for a ratio of two variables, so that's possible.. The way I'm most used to and I think what most people who know a little algebra are used to is to set up a system of linear equations. i.e. list the 5 unknowns: A, D, E, N, S Then list the equations in a consistent way, for example: 6A-S=0 0.4A-D=0 0.8E-N=0 A-D-E=o

OpenStudy (anonymous):

If you can rearrange and substitute things to get a new equation that contains only N and S as unknowns, then you can find the ratio, N/S.

OpenStudy (anonymous):

I'm going to try and give it a shot...Question what part of algebra is this?

OpenStudy (anonymous):

Basic equation solving, I guess. Systems of linear equations, more specifically.

OpenStudy (anonymous):

These questions are for my Finance class and the Professor says these variables will mean something later on in the semester...This is suppose to be a review of algebra and I dropped out of school when I was 17... I turn 34 in November.

OpenStudy (anonymous):

Ah, I see. Yes, this will definitely help you with the fundamentals that you will need for financial equations. Here's a hint: try to solve for A, D, and E in terms of only S or N, then put those into the A=D+E equation and you'll have all your S's and N's together to rearrange and find their ratio.

OpenStudy (anonymous):

I got my GED at 32 and returned to school...I passed college algebra my first semester in 2009 and haven't taken a math class since...Do you know of any sites that I can go to try and study.

OpenStudy (anonymous):

That A=D+E equation does look like something familiar from financial accounting. Though, I saw it as: Assets = Liabilities + Equity.

OpenStudy (anonymous):

There are a bunch of good ones. This one (Openstudy) is wonderful, probably the best for real-time interactive help. the Khan Academy videos on Youtube are great. My recommended sites are www.themathpage.com and do a search for "Paul's Online Math Notes" they are hosted at lamar.edu.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

n.p. good luck. If you find the solution, please post it so I can check my own work.

OpenStudy (anonymous):

ok